# Polarizable Water Model with Ab Initio Neural Network Dynamic Charges and Spontaneous Charge Transfer

### v1

@@ 3.3 Demonstrative application
As an illustrative application, we probe the origin of the emerging on-water reactivity on the air/water interface using ChargeNN. The electronic process for creating the strong electric fields on the water microdroplet surface remains elusive, and has been attributed to the abundance of dangling OH bonds at the interface 138, 139, the presence of OH-/H+ due to proton transfer 140, 141, 142 and electron transfer 143, 14, 15. To understand the source of the interfacial electric fields, we carried out statistical analysis to identify differences in geometries and charge distributions between the droplet and bulk liquid. We extracted the last 100 structural snapshots at 100 femtosecond intervals from a 1-nanosecond (ns) MD trajectory of a large water droplet with a radius $\sim$45 Å containing 10327 water molecules. For the liquid simulation, we analyzed the last 500 snapshots from a 1-ns MD trajectory of a large unit cell containing 2123 water molecules.
Our results reveal that the primary factor influencing the layer charge density is the hydrogen-to-oxygen ratio. The volume charge density (VCD) has been extensively used as an indicator for the strength of the layer electric fields143, and is highly correlated with the proportion of the hydrogen atoms (PHA) (see Figure 4b). For deep layers inside the droplet at distances less than 37 Å from the center, the PHA fluctuates around an equilibrium percentage of 67% indicating an homogeneous distribution, with the VCD oscillating near the liquid VCD of 0 e/nm3. The PHA on the intermediate layer at the distance of 41 Å from the center drops significantly to 66.2% , and coincides with the most negatively charged layer with a VCD of -0.35 e/nm3. Beyond this layer, the PHA grows continuously to break the theoretical limit of 66.7% (that is, 2 H atoms and 1 O atom per water) and even exceed 68% at the air/water interface, revealing an inhomogeneous distribution of H-bond network near the surface, where the VCD increases consistently and reaches a maximum of 0.18 e/nm3 at 45 Å. Our results align with the ”dangling OH theory” 138, 139 that more OH bonds dangle on the droplet surface due to the partial collapse of the hydrogen-bond network near the surface, leading to a higher proportion of hydrogen atoms and a positively charged surface layer. Consequently, the inward layer near the interface is abundant in oxygen atoms, and overall negatively charged.
@@ 4 Conclusions
Additionally, we probed the origin of the strong local electric fields on the droplet surface by conducting molecular dynamics of both liquid water and a large droplet using the ChargeNN water model. The findings clearly indicate that the layer electric fields on the droplet surface are primarily induced by breakage of the hydrogen-bond network, which leads to distinct proportion of hydrogen atoms from the liquid phase. Furthermore, surface-to-interior charge transfers also considerably contribute to the layer electron density.
Marching on, we can envisage broad applications with ChargeNN. Firstly, integrating this water model with QM electron structures via a QM/MM interface will allow accurate descriptions of outer solvent shells for reactive centers. Additionally, the ChargeNN-based water model can also be seamlessly combined with the existing force fields to simulate interactions between proteins and water solvents. Such interactions markedly affect the structures and dynamic behavior of proteins144. Moreover, incorporating empirical valence bond (EVB)145 and reactive force fields (ReaxFF)131, 132 into ChargeNN will enable simulations of bond breaking and formation, facilitating efficient modeling of proton transfers through hydrogen-bond networks and chemical reactions. A recent study shows that QM electronic polarization is essential for accurately producing spectral densities for proteins.146 Therefore, we aim to extend the ChargeNN to other atoms and molecular systems, enabling generalizable predictions of atomic charges across diverse local environments and achieving QM-level charge assignments for complex macromolecules. Combining accurate charges and force fields will allow for the depiction of instantaneous polarization and transient charge transfer in molecular dynamics. Coupled with well-designed energy functionals, fast prediction of ab initio polarization from neural networks would significantly enhance the accuracy of atomistic simulations for large chemical systems beyond water..
@@
The atomic charges were calculated using the third-order many-body-expansion orbital-specific-virtual MP2 (MBE(3)-OSV-MP2) method. The OSV approximation has demonstrated a significant reduction in the computational effort required for correlated methods without considerable loss in accuracy.147, 148, 149, 150, 112, 151, 152 The OSVs $(\mathbf{Q}_{k})$ can be obtained through a simple singular value decomposition of the diagonal pair MP2 amplitudes $\mathbf{T}_{kk}$:
$[\mathbf{Q}_{k}^{\dagger}\mathbf{T}_{kk}\mathbf{Q}_{k}]_{\overline{\mu}_{k}\overline{\nu}_{k}}=\omega_{\overline{\mu}_{k}}\delta_{\overline{\mu}\overline{\nu}}$ (7)
@@
where $\mathbf{T}_{ii}$ denotes the diagonal semi-canonical amplitudes, while $\mathbf{X}^{\mathrm{T}}_{ij}$ and $\mathbf{X}^{\bot}_{ji}$ represent the upper and lower diagonal blocks of the matrix containing OSV-MP2 responses.150
Additionally, we developed a large-scale parallelism strategy utilizing efficient memory management and data communication, which enables computations of MP2-level charges for large molecular systems.112
Charge Model 5
Charge Model 5 (CM5)26 based on Hirshfeld population analysis (HPA)25 was implemented for MBE(3)-OSV-MP2 to investigate the water polarization and charge transfers in response to the variation in the local environment. The Hirshfeld atomic charges are obtained by integrating the electron density within the region surrounding each atom,
$q^{\mathrm{HPA}}_{i}=Z_{i}-\int\mathrm{d}\mathbf{r}\frac{\rho^{0}_{i}(\mathbf{r})}{\sum_{j}\rho^{0}_{j}(\mathbf{r})}\rho(\mathbf{r})$ (11)
where $Z_{i}$ is the nuclear charge. $\rho^{0}(\mathbf{r})$ and $\rho(\mathbf{r})$ denote the electron densities of the promolecule and real molecule at position $\mathbf{r}$, respectively. The Hirshfeld population analysis provides a clear partitioning of the electron density and accurately describes the density redistribution from promolecule to real molecule during bonding. The CM5 model introduces a unified set of parameters to adjust the underestimated Hirshfeld charges, providing a more accurate description of the electrostatic potential.
$q^{\mathrm{CM5}}_{i}=q^{\mathrm{HPA}}_{i}+\sum_{j\neq i}A_{ij}B_{ij}$ (12)
where $A_{ij}$ is a coefficient and $B_{ij}$ refers to Pauling bond order. CM5 has been shown to deliver precise electronic charge distributions and be less sensitive to different theoretical levels and basis sets.26, 27. The CM5 charges in this work were obtained with $q^{\mathrm{HPA}}_{i}$, which was computed from the MBE(3)-OSV-MP2/cc-pVTZ relaxed density matrix. {acknowledgement}
The authors acknowledge financial supports from the Hong Kong Research Grant Council (17309020, 17305724), the Hong Kong Quantum AI Lab through the RTH-ITF program of the Hong Kong Government, and the Hung Hing Ying Physical Sciences Research Fund of the University of Hong Kong. Q.L. acknowledges Xinyan Wang and Ruiyi Zhou for valuable discussions.
@@
The Supporting Information is available free of charge.

### v3

@@ 3.3 Demonstrative application
As an illustrative application, we probe the origin of the emerging on-water reactivity on the air/water interface using ChargeNN. The electronic process for creating the strong electric fields on the water microdroplet surface remains elusive, and has been attributed to the abundance of dangling OH bonds at the interface 156, 157, the presence of OH-/H+ due to proton transfer 158, 159, 160 and electron transfer 161, 14, 15. To understand the source of the interfacial electric fields, we carried out statistical analysis to identify differences in geometries and charge distributions between the droplet and bulk liquid. We extracted the last 100 structural snapshots at 100 femtosecond intervals from a 1-nanosecond (ns) MD trajectory of a large water droplet with a radius $\sim$45 Å containing 10327 water molecules. For the liquid simulation, we analyzed the last 500 snapshots from a 1-ns MD trajectory of a large unit cell containing 2123 water molecules.
Our results reveal that the primary factor influencing the layer charge density is the hydrogen-to-oxygen ratio. The volume charge density (VCD) has been extensively used as an indicator for the strength of the layer electric fields161, and is highly correlated with the proportion of the hydrogen atoms (PHA) (see Figure 6b). For deep layers inside the droplet at distances less than 37 Å from the center, the PHA fluctuates around an equilibrium percentage of 67% indicating an homogeneous distribution, with the VCD oscillating near the liquid VCD of 0 e/nm3. The PHA on the intermediate layer at the distance of 41 Å from the center drops significantly to 66.2% , and coincides with the most negatively charged layer with a VCD of -0.35 e/nm3. Beyond this layer, the PHA grows continuously to break the theoretical limit of 66.7% (that is, 2 H atoms and 1 O atom per water) and even exceed 68% at the air/water interface, revealing an inhomogeneous distribution of H-bond network near the surface, where the VCD increases consistently and reaches a maximum of 0.18 e/nm3 at 45 Å. Our results align with the “dangling OH theory” 156, 157 that more OH bonds dangle on the droplet surface due to the partial collapse of the hydrogen-bond network near the surface, leading to a higher proportion of hydrogen atoms and a positively charged surface layer. Consequently, the inward layer near the interface is abundant in oxygen atoms, and overall negatively charged.
@@
4 Application Outlook
Beyond the simulation of water, we can envisage broader applications with ChargeNN. Firstly, the ChargeNN water model can be seamlessly combined with the existing force fields to simulate interactions between proteins and water solvents. Such interactions markedly affect the structures and dynamic behavior of proteins162. Additionally, integrating chargeNN water model with QM electron structures via a QM/MM interface will allow accurate descriptions of outer solvent shells for reactive centers. Moreover, incorporating empirical valence bond (EVB)163 and reactive force fields (ReaxFF)150, 151 into ChargeNN will enable simulations of bond breaking and formation, facilitating efficient modeling of proton transfers through hydrogen-bond networks and chemical reactions. Furthermore, accurate simulation of protein-ligand electrostatic interactions requires incorporating electronic polarizability, which induces screening effects that weaken electrostatics in the buried environment.164 A recent study also shows that QM electronic polarization is essential for accurately producing spectral densities for proteins.165 Therefore, we aim to extend the ChargeNN to other atoms and molecular systems, enabling generalizable predictions of atomic charges across diverse local environments and achieving QM-level charge assignments for complex macromolecules. Combining accurate charges and force fields will allow for the depiction of instantaneous polarization and transient charge transfer in molecular dynamics. Coupled with well-designed energy functionals, fast prediction of ab initio polarization from neural networks would significantly enhance the accuracy of atomistic simulations for large chemical systems other than water.
@@ 5 Conclusions
Additionally, we probed the origin of the strong local electric fields on the droplet surface by conducting molecular dynamics of both liquid water and a large droplet using the ChargeNN water model. The findings clearly indicate that the layer electric fields on the droplet surface are primarily induced by breakage of the hydrogen-bond network, which leads to distinct proportion of hydrogen atoms from the liquid phase. Furthermore, surface-to-interior charge transfer also considerably contributes to the layer electron density.
@@
The atomic charges were calculated using the MBE(3)-OSV-MP2 method. The OSV approximation has demonstrated a significant reduction in the computational effort required for correlated methods without considerable loss in accuracy.166, 167, 168, 169, 129, 170, 171 The OSVs $(\mathbf{Q}_{k})$ can be obtained through a simple singular value decomposition of the diagonal pair MP2 amplitudes $\mathbf{T}_{kk}$:
$[\mathbf{Q}_{k}^{\dagger}\mathbf{T}_{kk}\mathbf{Q}_{k}]_{\overline{\mu}_{k}\overline{\nu}_{k}}=\omega_{\overline{\mu}_{k}}\delta_{\overline{\mu}\overline{\nu}}.$ (11)
@@
where $\mathbf{T}_{ii}$ denotes the diagonal semi-canonical amplitudes, while $\mathbf{X}^{\mathrm{T}}_{ij}$ and $\mathbf{X}^{\bot}_{ji}$ represent the upper and lower diagonal blocks of the matrix containing OSV-MP2 responses.169
Additionally, we developed a large-scale parallelism strategy utilizing efficient memory management and data communication, which enables computations of MP2-level charges for large molecular systems.129
{acknowledgement}
The authors acknowledge financial supports from the Hong Kong Research Grant Council (17309020, 17305724), Hong Kong Quantum AI Lab through AIR@InnoHK program of Hong Kong Government, and the Hung Hing Ying Physical Sciences Research Fund of the University of Hong Kong. Q.L. acknowledges Dr. Xinyan Wang and Dr. Ruiyi Zhou for valuable discussions.
@@
The Supporting Information is available free of charge. The SI document comprises the results of the convergence of atomic charges and charge transfer with respect to atomic distances, testing MAEs relative to the training set size, parity graphs of molecular charges and charge transfer, optimized parameters for the ChargeNN water model, computational details for benchmarking ChargeNN, and some supplementary benchmark results.